OpenStudy (anonymous):

find the x and y intercept of the line that passes through the given points (-2,8),(4,2)

OpenStudy (anonymous):

i know the slope which is -6/6 0r -1

OpenStudy (anonymous):

then i got y-8=-1(x+2)

OpenStudy (anonymous):

what do i do next?

OpenStudy (wyattp17):

oh thats easy use the y intercept formula y = mx + b AND JUST plug in each

OpenStudy (wyattp17):

im i correct?

OpenStudy (anonymous):

but what will b be will it be x1 or y1

OpenStudy (haseeb96):

m=y2-y1\x2-x1 (x1,y1)=(-2,8) (x2,y2)=(4,2) m=2-8\4+2 m=-6\6 m=-1 by using point slope form y=mx+b y=-x+b for x intercepts put y=0 0=-x+b x=b for y intercapt put x=0 y=b both have same intercept points

OpenStudy (anonymous):

oooohhhhhhh

OpenStudy (wyattp17):

Can i plz have a medal for trying :) and giving u the formula?

OpenStudy (wyattp17):

@angelkitty456

OpenStudy (anonymous):

also will why be -x then for y or -1

OpenStudy (wyattp17):

yes -x would just be -1 and I'm not sure about y

OpenStudy (anonymous):

the answer is x is 6 and y is 6 but how do i show my work

OpenStudy (wyattp17):

hmmm @angelkitty456

OpenStudy (anonymous):

(3,-8),(-4,13)

OpenStudy (jonnyvonny):

Are you trying to find the slope?

OpenStudy (anonymous):

no im trying to find the and y intercepts

OpenStudy (anonymous):

the x and y intercepts

OpenStudy (anonymous):

you can find the x and y intercepts from the slope and the points you are given

OpenStudy (anonymous):

well for x i got 2 and for y i got 3 i think that is wrong though

OpenStudy (jonnyvonny):

Ok, so, we first need to find the slope of this function: use the slope formula: (y2-y1)/(x2-x1)=m\[(13-(-8))/(-4-3)= (13+8)/(-7)=21/-7=-3\] From then, you know that the slope, m=-3. From there, you are able to find b, the y-intercept with the formula y=mx+b; you know y, x, and m, just solve for b. The way you do this, is plug in any of the two points given, then slove:\[13=(-2)(-4)+b \rightarrow13=8+b \rightarrow 5=b\]With us knowing now the y-intercept, we are enabled to apply the slope for ANY point: y=mx+b... y=-2x+5 Since we found the y-intercept, b , which was 5, we just need to find the x-intercept; so make the function equal to 0.\[y=-2x+5\rightarrow 0=-2x+5\rightarrow-5=-2x \rightarrow(5/2)=x\] Hope this helped, if you need more elaboration, feel free to type back :) @angelkitty456

OpenStudy (anonymous):

i have another it is (5,0.4)

OpenStudy (anonymous):

(5,0.4) AND (-1,-2)

OpenStudy (jonnyvonny):

You should practice by solving it, and checking your answer (I will check it, if you so desire).

OpenStudy (anonymous):

i found thatand i got it but i had a new set posted above (5,0.4),(-1,-2) please help me

OpenStudy (anonymous):

for slope i got -2,4/6 but then what do i do